A married couple is planning to have 2 children and is interested in knowing the probabilities of having 0 girls, 1 girl or 2 girls. Assume the probability of having a girl is 0.50 in any one birth.
construct a probability distribution of the random variable x, where x is the # of girls born to this couple..
Thanks

Answer by stanbon(48568)

P(x=0)=3C0 1/2)^0 * 1/2)^3=1/8
P(x=1)=3C1 (1/2)^1 * (1/2)^2 = 3*1/8 = 3/8
P(x=2)=3C2 (1/2)^2 * 1/2)^1 = 3*1/8 = 3/8
P(x=3)=3C3 (1/2)^3 * (1/2)^0 = 1*(1/2)^3*(1/2)^0 = 1/8
-------
Cheers,
Stan H.